OpenStudy (anonymous):

Perform the indicated operation -5/9 - (-2/3).. thanks in advance

OpenStudy (anonymous):

since your minusing a negtive number the sign will change. -5/9+2/3 then. get the least common multiple.which is 9. -5/9 +2/3 = -5+6/9 -5/9+2/3=1/9 the answer is 1/9 .

OpenStudy (anonymous):

how did you get the -5+6/9

OpenStudy (anonymous):

having the least common multiple(LCM) as the denominator divide 9 from the denominator of the first fraction which is 9 then multiply it to the numerator which is -5. to get +6...same procedure. :D

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

try it. and see if you get the same answer.:)